# Security Code Review by Large Language Models

### v2

@@ 3.5.2. Regression Model Construction
Correlation and Redundancy Analysis: If explanatory variables are highly correlated with each other, it can cause interference in model construction and analysis. As the predefined explanatory variables in our study include both categorical and continuous variables, we utilized a refined Pearson’s hypothesis test of independence — Phi_K (Baak et al., 2020), to measure the correlation between each factor. The correlation coefficients obtained are illustrated in Fig. 2. We chose $|\rho|>0.7$ as the threshold because it is recommended as the threshold for strong correlation (Hinkle et al., 2003). From Fig. 2, where each factor is denoted by an abbreviation, it can be seen that the correlation coefficients between each pair of factors are all below 0.7, suggesting that there is no strong correlation between these factors. However, two factors without high correlation can still be duplicates, which can distort the relationship between explanatory and response variables in model fitting. We encoded categorical variables as dummy variables for redundancy analysis (Legendre and Legendre, 2012). Utilizing the redun function in the rms package, we assessed potential redundant variables with its default threshold of $R^{2}\geq 0.9$, ultimately finding none.
Figure 2. The Phi_K correlation coefficients of explanatory variables.
@@
Degrees of Freedom Allocation: To assign degrees of freedom to each explanatory variable, we measured the potential of nonlinear relationships between each factor and the response variable by calculating the Spearman rank correlations (${\rho}^{2}$) between them. As recommended by Harrell  (Harrell, 2001), explanatory variables with larger ${\rho}^{2}$ can be assigned more degrees of freedom within the range of 3 to 5. According to the computation results shown in Fig. 3, for interval variables, we considered using the rcs function from the rms package to allocate 3 degrees of freedom for the factor ‘TK’ with the highest ${\rho}^{2}$, while the other factors are allocated 1 degree of freedom. For categorical variables, they should be converted into factor variables, and their corresponding degrees of freedom are automatically allocated during model fitting.
@@
Figure 3. Dotplot of the Spearman ${\rho}^{2}$ of explanatory variable and response variable in three repetitive experiments
@@
Model Fitting: Next, we fitted the 1,215 responses and their rating categories onto a cumulative link mixed model using the clmm function from the ordinal R package (Christensen and Christensen, 2015). To balance model complexity and fit, we tested models with and without allocating 3 degrees of freedom to the ‘TK’ factor. The model that did not allocate these degrees of freedom, which had a lower Akaike Information Criterion (AIC), was then selected as our final fitting result (Burnham and Anderson, 2004).
@@
All the results and scripts utilized in data labelling, extraction, and regression analysis are provided in our replication package (Anonymous, 2024).
4. Results and Analysis
4.1. RQ1 Performance of LLMs
@@
Based on the evaluation metrics defined in Section 3.3.3, we evaluated the capability of six popular LLMs in security code review across five different prompts and compared the performance of LLMs with baseline tools in the C/C++ and Python dataset, respectively. If two out of the three metrics were superior (i.e., higher I-Score or IH-Score, lower M-Score), we deemed that the current combination outperforms the other ones. Due to differences in the token limit of each LLM and the inherent token count of each prompt template, the number of responses successfully generated by each LLM-prompt combination also varied. To ensure rigorous comparisons, we selected the intersection of source code files in our complete dataset so that all LLM-prompt combinations could generate responses, resulting in 57 instances. As a result, two sets of scores are provided: 1) the response count and performance scores of each LLM-prompt combination on the complete dataset (Table 4), and 2) the performance scores of each LLM-prompt combination on the selected 57 instances (Table 5).
@@
From Table 4, it is evident that, on the Python dataset, the combination of GPT-4 with $\textbf{P}_{cid}$ achieved the best performance. However, on the C/C++ dataset, the combination of Llama 2 70B with $\textbf{P}_{cid}$ performs the best. However, Table 5 demonstrated that in the selected 57 cases, Llama 2 70B scored obviously lower than GPT-4. We speculate that this discrepancy could be attributed to the smaller volume of responses obtained by Llama 2 70B, leading to inflated performance scores of Llama 2 70B in Table 4. On the whole, we consider GPT-4 paired with $\textbf{P}_{cid}$ as the top-performing LLM-prompt combination. As of the time of our experiment (May 2024), the capabilities of currently popular LLMs in conducting security code review are still limited. Notably, with enhanced prompts, the best performance of each LLM in security code review is significantly superior to that of static analysis tools.
@@
Table 4. Performance scores and entropy for each LLM-Prompt combination on the python and c/c++ dataset
Python C/C++
Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$ Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
GPT-3.5 Resp. Count 123 123 123 123 122 GPT-3.5 Resp. Count 141 141 141 139 137
I-Score 1.08% 1.35% 0.27% 0.27% 1.37% I-Score 1.89% 2.13% 0.00% 0.24% 0.73%
IH-Score 2.71% 3.80% 1.35% 0.27% 5.19% IH-Score 5.67% 3.31% 0.47% 0.96% 1.46%
M-Score 89.37% 84.28% 92.41% 97.29% 69.12% M-Score 76.60% 86.05% 99.29% 95.44% 84.43%
Entropy 0.2572 0.3692 0.2070 0.0746 0.6842 Entropy 0.4860 0.2439 0.0195* 0.1153 0.2931
GPT-4 Resp. Count 175 175 175 174 174 GPT-4 Resp. Count 247 247 247 247 247
I-Score 4.76% 3.43% 3.24% 5.17*3.06% I-Score 4.45% 3.91% 3.91% 5.40% 5.26%
IH-Score 9.34% 9.52% 8.38% 14.56*8.62% IH-Score 7.96% 8.01% 9.31% 11.74% 11.47%
M-Score 61.31% 60.95% 46.47% 44.06% 61.69% M-Score 55.47% 64.24% 44.53% 39.00% 43.59%
Entropy 0.7426 0.7283 0.8600 0.9416 0.7450 Entropy 0.6215 0.5329 0.6505 0.6955 0.7330
GPT-4 Turbo Resp. Count 201 201 201 201 201 GPT-4 Turbo Resp. Count 269 269 269 269 269
I-Score 0.66% 0.33% 0.66 0.83 0.17 I-Score 0.25% 0.12% 0.25% 0.50% 0.25%
IH-Score 0.66 0.83 2.32 1.99 0.33 IH-Score 0.25% 0.37% 1.24% 1.61% 1.12%
M-Score 97.18% 96.35% 85.74% 92.87% 98.01% M-Score 98.02% 98.64% 91.08% 93.43% 96.28%
Entropy 0.1757 0.0746 0.3370 0.6125 0.0404* Entropy 0.1430 0.0457 0.2493 0.5413 0.0250
Gemini Pro Resp. Count 162 161 160 160 159 Gemini Pro Resp. Count 223 221 221 221 220
I-Score 0.00% 0.00% 0.63% 1.25% 0.21% I-Score 0.00% 0.15% 0.30% 1.66% 0.00%
IH-Score 0.41% 0.00% 0.84% 3.54% 0.21% IH-Score 0.45% 0.15% 0.90% 3.77% 0.45%
M-Score 90.54% 97.09% 84.04% 55.21% 98.53% M-Score 92.97% 98.19% 89.44% 61.54% 99.09%
Entropy 0.0777 0.0901 0.2424 0.1873 0.0548 Entropy 0.0512 0.0332 0.1401 0.1576 0.1024
Llama 2 7B Resp. Count 62 62 61 61 50 Llama 2 7B Resp. Count 57 57 57 51 33
I-Score 0.00% 0.54% 1.62% 2.19% 1.33 I-Score 0.58% 0.00% 0.00% 2.61% 1.01%
IH-Score 1.61% 1.61% 5.40% 8.74% 4.67% IH-Score 2.34% 3.51% 5.85% 4.58% 7.07%
M-Score 87.10% 83.33% 56.78% 40.98% 33.33*M-Score 90.06% 87.13% 81.29% 52.94% 27.27*%
Entropy 0.2625 0.3070 0.4766 0.4351 0.5521 Entropy 0.2650 0.2167 0.2167 0.4092 0.5954
Llama 2 70B Resp. Count 62 62 62 61 47 Llama 2 70B Resp. Count 57 57 57 51 37
I-Score 0.54% 1.07% 2.15% 3.83% 1.42% I-Score 1.17% 1.17% 1.75% 6.54*1.80%
IH-Score 4.84% 4.30% 6.45% 11.48% 3.55% IH-Score 4.68% 2.92% 3.51% 15.03*6.31%
M-Score 79.57% 84.41% 58.60% 43.17% 36.17% M-Score 67.25% 80.12% 66.08% 40.52% 54.05%
Entropy 0.4470 0.3474 0.3514 0.3832 0.5762 Entropy 0.5667 0.4145 0.3895 0.5515 0.6153
Baselines SAST SonarQube CodeQL Bandit Semgrep Baselines SAST CppCheck Semgrep
Resp. Count 258 258 258 258 Resp. Count 276 276
I-Score 0.00% 0.39% 0.00% 0.00% I-Score 0.00% 0.00%
IH-Score 0.00% 1.16% 1.16% 0.00% IH-Score 0.72% 0.36%
M-Score 91.47% 82.56% 78.29% 90.31% M-Score 98.19% 96.74%
  • The best values for each score are highlighted in bold and marked with *.
@@
When using basic prompts constructed by $\textbf{P}_{b}$ (see Table 4), GPT-4 performs the best, obtaining the highest I-Score (Python: 4.76%, C/C++: 4.45%), IH-Score (Python: 9.34%, C/C++: 7.96%), and the lowest M-Score (Python: 61.31%, C/C++: 55.47%). Following closely behind is Llama 2 70B and GPT-3.5. GPT-4 Turbo, Gemini Pro, and Llama 2 7B performed the worst, with their performance in the Python dataset even inferior to that of the baseline tool, Bandit. Similar results are observed in Table 5. Specifically, in the selected 57 instances, under the basic prompt $\textbf{P}_{b}$, GPT-4 achieved I-Score = 8.77%, IH-Score = 20.46%, and M-Score = 43.86%. It is evident that GPT-4’s performance on the selected 57 instances exceeds its performance across the entire dataset. Given that these 57 code files are instances with relatively fewer tokens within our dataset, it could be speculated that the quantity of tokens input to LLM affects its performance in detecting security defects, which has been further explored and analyzed in RQ3.
@@
As shown in Table 4, the changes of the performance scores of each LLM under prompt $\textbf{P}_{r}$ compared with that under $\textbf{P}_{b}$ is generally small. Specifically, GPT-3.5, GPT-4, and GPT-4 Turbo show subtle increases and decreases in performance on the Python and C/C++ datasets, respectively. Gemini Pro and Llama 2 70B demonstrate minor performance decreases on both datasets, while Llama 2 7B shows a slight improvement. A possible reason is that the name of components or packages utilized may already imply the source project of the code file, thereby diminishing the impact of project information provided in $\textbf{P}_{r}$. Furthermore, the improvement of emphasizing the task through adding a persona in the prompt is LLM-specific. We speculate that some LLMs may already understand the task described in the prompt quite well, thus introducing a persona could add unnecessary complexity of the prompt and lead to distractions.
We can see from Table 4 that compared to using the basic prompt $\textbf{P}_{b}$, the performance of GPT-3.5 significantly decreases under $\textbf{P}_{c}$ and $\textbf{P}_{cid}$ in both Python and C/C++ datasets. The performances of GPT-4 with $\textbf{P}_{c}$ exhibit different changes across different datasets, with an improvement in the C/C++ dataset but a decrease in the Python dataset. However, all other LLMs achieved higher performance scores across both datasets. Moreover, GPT-4, GPT-4 Turbo, Gemini Pro, Llama 2 7B, and Llama 2 70B all achieved their best performance under $\textbf{P}_{cid}$. Therefore, it can be inferred that adding information related to CWE to the prompt generally enhances the capabilities of security defect detection of various LLMs in code review. It is worth noting that under $\textbf{P}_{cid}$, the decrease in GPT-4’s I-Score (Python: 4.76% →3.24%, C/C++: 4.45% →3.91%) and in M-Score (Python: 61.31% →46.47%, C/C++:55.47% →44.53%) reflect that guiding LLMs to reference CWE in the prompt indeed enables GPT-4 to detect more security defects, but it does not guarantee the accuracy of the detection results. Under $\textbf{P}_{cid}$, providing a specific CWE list as a reference in the prompt leads to significant improvements across all metrics and datasets for GPT-4. Hence, how to leverage CWE information in the prompt is a question worthy of consideration.
@@
Table 4 illustrates that $\textbf{P}_{cot}$ significantly enhances the ability to detect security defects for Llama 2 7B and Llama 2 70B. However, for Gemini Pro and GPT-4 Turbo, there is no improvement and even a detrimental effect to these two LLMs. As for GPT-3.5 and GPT-4, they respectively exhibit improvement and decrease in their performances across different datasets. Such instability may be attributed to differences in training data among different LLMs. Additionally, we first instructed the LLM to generate missing code context based on the corresponding commit message, where the consistency between this generated context and the actual code context cannot be guaranteed. Through analyzing specific cases, we discovered that some LLMs automatically fixed security defects while generating code context, leading to confusing and erroneous detection results in the second step of detecting security defects in the provided code.
@@
According to the mean entropies of various LLM-prompt combinations in Table 4, GPT-3.5, GPT-4 Turbo, and Gemini Pro exhibit relatively higher consistency in their responses across three repetitive experiments, followed by Llama 2 7B and Llama 2 70B, with GPT-4 showing the least consistency. Specifically, on the Python dataset, the highest consistency of responses was observed under $\textbf{P}_{cot}$ with GPT-4 Turbo, with a mean entropy of 0.0404. On the C/C++ dataset, the combination of GPT-3.5 and $\textbf{P}_{c}$ demonstrates the lowest mean entropy of 0.0195. Given the poor security defect detection capabilities of these two LLM-prompt combinations, their high consistency may be due to their tendency to not detect any security defects in the code. GPT-4 paired with $\textbf{P}_{cid}$ is considered the top-performing combination in security code review, but shows high mean entropies on both Python and C/C++ datasets, with 0.9416 and 0.6955 respectively. These results suggest that LLMs with stronger security defect detection capabilities may also exhibit greater randomness.
4.2. RQ2 Quality Problems in Responses
The average distribution of 12 problem types across 4 themes in responses to three repetitive experiments is illustrated in Fig. 4, where the same color represents types within the same theme, with darker shades indicating the themes.

### v3

@@ 3.5.2. Regression Model Construction
Correlation and Redundancy Analysis: If explanatory variables are highly correlated, it can cause interference in model construction and analysis. As the predefined explanatory variables in our study include both categorical and continuous variables, we utilized a refined Pearson’s hypothesis test of independence — Phi_K (Baak et al., 2020), to measure the correlation between each factor. The correlation coefficients obtained are illustrated in Fig. 3. We chose $|\rho|>0.7$ as the threshold because it is recommended as the threshold for strong correlation (Hinkle et al., 2003). From Fig. 3, where an abbreviation denotes each factor, it can be seen that the correlation coefficients between each pair of factors are all below 0.7, suggesting that there is no strong correlation between these factors. However, two factors without high correlation can still be duplicates, distorting the relationship between explanatory and response variables in model fitting. We encoded categorical variables as dummy variables for redundancy analysis (Legendre and Legendre, 2012). Utilizing the redun function in the rms package, we assessed potential redundant variables with its default threshold of $R^{2}\geq 0.9$, ultimately finding none.
Figure 3. The Phi_K correlation coefficients of explanatory variables
@@
Degrees of Freedom Allocation: To assign degrees of freedom to each explanatory variable, we measured the potential of nonlinear relationships between each factor and the response variable by calculating the Spearman’s rank correlations (${\rho}^{2}$) between them. As Harrell  (Harrell, 2001) recommended, explanatory variables with larger ${\rho}^{2}$ can be assigned more degrees of freedom within the range of 3 to 5. According to the computation results shown in Fig. 4, for interval variables, we considered using the rcs function from the rms package to allocate 3 degrees of freedom for the factor ‘Token’ with the highest ${\rho}^{2}$, while the other factors are allocated 1 degree of freedom. Category variables should be converted into factor variables, and their corresponding degrees of freedom are automatically allocated during model fitting.
@@
Figure 4. Dotplot of the Spearman ${\rho}^{2}$ between explanatory and response variables across three repetitive experiments
@@
Model Fitting: Next, we fitted the 1,215 responses and their rating categories onto a cumulative link mixed model using the clmm function from the ordinal R package (Christensen and Christensen, 2015). We tested models with and without allocating 3 degrees of freedom to the ‘Token’ factor to balance model complexity and fit. The model that did not allocate these degrees of freedom, which had a lower Akaike Information Criterion (AIC), was then selected as our final fitting result (Burnham and Anderson, 2004).
@@
All the results and scripts utilized in data labelling, extraction, and regression analysis are provided in our replication package (Yu et al., 2024).
4. Results
4.1. Performance of LLMs (RQ1)
@@
Based on the evaluation metrics defined in Section 3.3.3, we evaluated the capability of six popular LLMs in security code review across five different prompts. We compared the performance of LLMs with baseline tools in the C/C++ and Python datasets, respectively. If two of the three metrics were superior (i.e., higher I-Score or IH-Score, lower M-Score), we deemed that the current combination outperforms the others. Due to differences in the token limit of each LLM and the inherent token count of each prompt template, the number of responses successfully generated by each LLM-prompt combination also varied. To ensure rigorous comparisons, we selected the intersection of source code files in our complete dataset so that all LLM-prompt combinations could generate responses, resulting in 57 instances. As a result, two sets of scores are provided: (1) the response count and performance scores of each LLM-prompt combination on the complete Python and C/C++ dataset (see Table 4 and Table 5), and (2) the performance scores of each LLM-prompt combination on the selected 57 instances (see Table 6). The best values for each score are highlighted in bold and marked with asterisks in the tables.
@@
From Table 4 and Table 5, it is evident that, on the Python dataset, the combination of GPT-4 with $\textbf{P}_{cid}$ achieved the best performance. However, on the C/C++ dataset, the combination of Llama 2 70B with $\textbf{P}_{cid}$ performs the best. However, Table 6 demonstrated that in the selected 57 cases, Llama 2 70B scored lower than GPT-4. We speculate that this discrepancy could be attributed to the smaller volume of responses obtained by Llama 2 70B, leading to inflated performance scores of Llama 2 70B in the complete datasets. On the whole, we consider GPT-4 paired with $\textbf{P}_{cid}$ as the top-performing LLM-prompt combination. As of the time of our experiment (May 2024), the capabilities of currently popular LLMs in conducting security code reviews are still limited. Notably, with enhanced prompts, the best performance of each LLM in security code review is significantly superior to that of static analysis tools. This could be attributed to two reasons: (1) for context-sensitive tools such as CodeQL, static analysis can be impacted by the lack of code context in our dataset (Russo and Sabelfeld, 2010); (2) there is a relatively high frequency of security defects related to multi-threading and asynchronous programming, such as race conditions. As such features can only be accurately modelled using dynamic analysis, these security defects are challenging for static analysis tools to detect (Charoenwet et al., [n. d.]).
@@
Table 4. Performance scores and entropy for each LLM-Prompt combination on the Python dataset
Python
GPT-3.5 Resp. Count 123 123 123 123 122
I-Score 1.08% 1.35% 0.27% 0.27% 1.37%
IH-Score 2.71% 3.80% 1.35% 0.27% 5.19%
M-Score 89.37% 84.28% 92.41% 97.29% 69.12%
Entropy 0.2572 0.3692 0.2070 0.0746 0.6842
GPT-4 Resp. Count 175 175 175 174 174
I-Score 4.76% 3.43% 3.24% 5.17*3.06%
IH-Score 9.34% 9.52% 8.38% 14.56*8.62%
M-Score 61.31% 60.95% 46.47% 44.06% 61.69%
Entropy 0.7426 0.7283 0.8600 0.9416 0.7450
GPT-4 Turbo Resp. Count 201 201 201 201 201
I-Score 0.66% 0.33% 0.66 0.83 0.17
IH-Score 0.66 0.83 2.32 1.99 0.33
M-Score 97.18% 96.35% 85.74% 92.87% 98.01%
Entropy 0.1757 0.0746 0.3370 0.6125 0.0404*
Gemini Pro Resp. Count 162 161 160 160 159
I-Score 0.00% 0.00% 0.63% 1.25% 0.21%
IH-Score 0.41% 0.00% 0.84% 3.54% 0.21%
M-Score 90.54% 97.09% 84.04% 55.21% 98.53%
Entropy 0.0777 0.0901 0.2424 0.1873 0.0548
Llama 2 7B Resp. Count 62 62 61 61 50
I-Score 0.00% 0.54% 1.62% 2.19% 1.33
IH-Score 1.61% 1.61% 5.40% 8.74% 4.67%
M-Score 87.10% 83.33% 56.78% 40.98% 33.33*%
Entropy 0.2625 0.3070 0.4766 0.4351 0.5521
Llama 2 70B Resp. Count 62 62 62 61 47
I-Score 0.54% 1.07% 2.15% 3.83% 1.42%
IH-Score 4.84% 4.30% 6.45% 11.48% 3.55%
M-Score 79.57% 84.41% 58.60% 43.17% 36.17%
Entropy 0.4470 0.3474 0.3514 0.3832 0.5762
Baselines SAST SonarQube CodeQL Bandit Semgrep
Resp. Count 258 258 258 258
I-Score 0.00% 0.39% 0.00% 0.00%
IH-Score 0.00% 1.16% 1.16% 0.00%
M-Score 91.47% 82.56% 78.29% 90.31%
Table 5. Performance scores and entropy for each LLM-Prompt combination on the C/C++ dataset
C/C++
GPT-3.5 Resp. Count 141 141 141 139 137
I-Score 1.89% 2.13% 0.00% 0.24% 0.73%
IH-Score 5.67% 3.31% 0.47% 0.96% 1.46%
M-Score 76.60% 86.05% 99.29% 95.44% 84.43%
Entropy 0.4860 0.2439 0.0195* 0.1153 0.2931
GPT-4 Resp. Count 247 247 247 247 247
I-Score 4.45% 3.91% 3.91% 5.40% 5.26%
IH-Score 7.96% 8.01% 9.31% 11.74% 11.47%
M-Score 55.47% 64.24% 44.53% 39.00% 43.59%
Entropy 0.6215 0.5329 0.6505 0.6955 0.7330
GPT-4 Turbo Resp. Count 269 269 269 269 269
I-Score 0.25% 0.12% 0.25% 0.50% 0.25%
IH-Score 0.25% 0.37% 1.24% 1.61% 1.12%
M-Score 98.02% 98.64% 91.08% 93.43% 96.28%
Entropy 0.1430 0.0457 0.2493 0.5413 0.0250
Gemini Pro Resp. Count 223 221 221 221 220
I-Score 0.00% 0.15% 0.30% 1.66% 0.00%
IH-Score 0.45% 0.15% 0.90% 3.77% 0.45%
M-Score 92.97% 98.19% 89.44% 61.54% 99.09%
Entropy 0.0512 0.0332 0.1401 0.1576 0.1024
Llama 2 7B Resp. Count 57 57 57 51 33
I-Score 0.58% 0.00% 0.00% 2.61% 1.01%
IH-Score 2.34% 3.51% 5.85% 4.58% 7.07%
M-Score 90.06% 87.13% 81.29% 52.94% 27.27*%
Entropy 0.2650 0.2167 0.2167 0.4092 0.5954
Llama 2 70B Resp. Count 57 57 57 51 37
I-Score 1.17% 1.17% 1.75% 6.54*1.80%
IH-Score 4.68% 2.92% 3.51% 15.03*6.31%
M-Score 67.25% 80.12% 66.08% 40.52% 54.05%
Entropy 0.5667 0.4145 0.3895 0.5515 0.6153
Baselines SAST CppCheck Semgrep
Resp. Count 276 276
I-Score 0.00% 0.00%
IH-Score 0.72% 0.36%
M-Score 98.19% 96.74%
@@
When using basic prompts constructed by $\textbf{P}_{b}$ (see Table 4 and Table 5), GPT-4 performs the best, obtaining the highest I-Score (Python: 4.76%, C/C++: 4.45%), IH-Score (Python: 9.34%, C/C++: 7.96%), and the lowest M-Score (Python: 61.31%, C/C++: 55.47%). Following closely behind is Llama 2 70B and GPT-3.5. GPT-4 Turbo, Gemini Pro, and Llama 2 7B performed the worst, with their performance in the Python dataset even inferior to that of the baseline tool, Bandit. Similar results are observed in Table 6. Specifically, in the selected 57 instances, under the basic prompt $\textbf{P}_{b}$, GPT-4 achieved I-Score = 8.77%, IH-Score = 20.46%, and M-Score = 43.86%. It is evident that GPT-4’s performance on the selected 57 instances exceeds its performance across the entire dataset. Given that these 57 code files are instances with relatively fewer tokens within our dataset, it could be speculated that the quantity of tokens input to LLM affects its performance in detecting security defects, which has been further explored and analyzed in RQ3.
@@
As shown in Table 4 and Table 5, the changes of the performance scores of each LLM under prompt $\textbf{P}_{r}$ compared with that under $\textbf{P}_{b}$ is generally small. Specifically, GPT-3.5, GPT-4, and GPT-4 Turbo show subtle increases and decreases in performance on the Python and C/C++ datasets, respectively. Gemini Pro and Llama 2 70B demonstrate minor performance decreases on both datasets, while Llama 2 7B shows a slight improvement. A possible reason is that the name of components or packages utilized may already imply the source project of the code file, thereby diminishing the impact of project information provided in $\textbf{P}_{r}$. Furthermore, the improvement of emphasizing the task through adding a persona in the prompt is LLM-specific. We speculate that some LLMs may already understand the task described in the prompt quite well, and thus, introducing a persona could add unnecessary complexity to the prompt and lead to distractions.
We can see from Table 4 and Table 5 that compared to using the basic prompt $\textbf{P}_{b}$, the performance of GPT-3.5 significantly decreases under $\textbf{P}_{c}$ and $\textbf{P}_{cid}$ in both Python and C/C++ datasets. The performances of GPT-4 with $\textbf{P}_{c}$ exhibit different changes across different datasets, with an improvement in the C/C++ dataset but a decrease in the Python dataset. However, all other LLMs achieved higher performance scores across both datasets. Moreover, GPT-4, GPT-4 Turbo, Gemini Pro, Llama 2 7B, and Llama 2 70B all achieved their best performance under $\textbf{P}_{cid}$. Therefore, it can be inferred that adding information related to CWE to the prompt generally enhances the capabilities of security defect detection of various LLMs in code review. It is worth noting that under $\textbf{P}_{cid}$, the decrease in GPT-4’s I-Score (Python: 4.76% →3.24%, C/C++: 4.45% →3.91%) and in M-Score (Python: 61.31% →46.47%, C/C++:55.47% →44.53%) reflect that guiding LLMs to reference CWE in the prompt indeed enables GPT-4 to detect more security defects. Still, it does not guarantee the accuracy of the detection results. Under $\textbf{P}_{cid}$, providing a specific CWE list as a reference in the prompt leads to significant improvements across all metrics and datasets for GPT-4. Hence, how to leverage CWE information in the prompt is worthy of consideration.
@@
Table 4 and Table 5 illustrate that $\textbf{P}_{cot}$ significantly enhances the ability to detect security defects for Llama 2 7B and Llama 2 70B in both Python and C/C++ datasets. However, for Gemini Pro and GPT-4 Turbo, there is no improvement and even a detrimental effect to these two LLMs. As for GPT-3.5 and GPT-4, they respectively exhibit improvement and decrease in their performances across different datasets. It can be seen that the impact of supplementing missing context by CoT prompting on the performance of LLMs is multifaceted and unstable. A possible explanation is that different LLMs vary in their ability to process and understand input. Moreover, we instructed the LLM to generate missing code context based on the corresponding commit message in $\textbf{P}_{cot}$. On the one hand, the generated context may contain the critical information needed to analyze the target security defect, which can help the LLM in security code review. On the other hand, the consistency between this generated context and the actual code context cannot be guaranteed. We observed that in some cases, the code context generated under Prompt $\textbf{P}_{cot-1}$ was inconsistent with the actual code context, resulting in the original security defect no longer being identified as a defect within the generated context. Consequently, when Prompt $\textbf{P}_{cot-2}$ was applied, the detection of security defects in the provided code was misled by the generated code context and yielded erroneous results.
@@
According to the mean entropies of various LLM-prompt combinations in Table 4 and Table 5, GPT-3.5, GPT-4 Turbo, and Gemini Pro exhibit relatively higher consistency in their responses across three repetitive experiments, followed by Llama 2 7B and Llama 2 70B, with GPT-4 showing the least consistency. Specifically, on the Python dataset, the highest consistency of responses was observed under $\textbf{P}_{cot}$ with GPT-4 Turbo, with a mean entropy of 0.0404. On the C/C++ dataset, the combination of GPT-3.5 and $\textbf{P}_{c}$ demonstrates the lowest mean entropy of 0.0195. Given these two LLM-prompt combinations’ poor security defect detection capabilities, their high consistency may be due to their tendency not to detect any security defects in the code. GPT-4 paired with $\textbf{P}_{cid}$ is considered the top-performing combination in security code review but shows high mean entropies on both Python and C/C++ datasets, with 0.9416 and 0.6955 respectively. These results suggest that LLMs with more robust security defect detection capabilities may also exhibit greater randomness.
4.2. Quality Problems in Responses (RQ2)
The average distribution of 12 problem types across four themes in responses to three repetitive experiments is illustrated in Fig. 5, where the same color represents types within the same theme, with darker shades indicating the themes.